### v2

Abstract
Remarkable strides in computational pathology (CPath) have been made in the task-agnostic foundation model (FM) that advances the performance of a wide array of downstream clinical tasks. Despite the promising performance, there are still several challenges. First, prior works have resorted to either vision-only or vision-captions data, disregarding invaluable pathology reports and gene expression profiles which respectively offer distinct knowledge for versatile clinical applications. Second, the current progress in pathology FMs predominantly concentrates on the patch level, where the restricted context of patch-level pretraining fails to capture whole-slide patterns. Here we curated the largest multimodal dataset consisting of H&E diagnostic whole slide images (WSIs) along with their associated pathology reports and RNA-Seq data, which resulted in 26,169 slide-level modality pairs from 10,275 patients across 32 cancer types. To leverage these data for CPath, we propose a novel whole-slide pretraining paradigm which injects multimodal knowledge at the whole-slide context into the pathology FM, called Multimodal Self-TAught PRetraining (mSTAR). The proposed paradigm revolutionizes the workflow of pretraining for CPath, which enables the pathology FM to acquire the whole-slide context. To the best of our knowledge, this is the first attempt to incorporate multimodal knowledge at the slide level for enhancing pathology FMs, expanding the modelling context from unimodal to multimodal knowledge and from patch-level to slide-level. To systematically evaluate the capabilities of mSTAR, extensive experiments including slide-level unimodal and multimodal applications, are conducted across 7 diverse types of tasks on 43 subtasks, resulting in the largest spectrum of downstream tasks. The average performance in varying types of slide-level applications consistently demonstrates significant performance enhancements for mSTAR compared to other state-of-the-art FMs, with statistically critical differences observed. In particular, mSTAR showcases the promising superiority in multimodal capabilities attained through the integration of multimodal knowledge.
keywords:
foundation models, computational pathology, multimodal, pretraining
1 Introduction
Figure 1: Overview of the study. a, The workflow in clinical practice for diagnosis, treatment and prognosis of oncology, which primarily involves three common modalities data: WSIs, pathology reports and gene expression profiles. b, The overview of mSTAR paradigm. mSTAR consists of two stages: 1) Slide-level Contrastive Learning, 2) Patch-level Self-Taught Training. c-e, statistics of data used in this study, including c) Venn Graph of cases across various modalities, d) the number of cases for different cancer types and e) the distribution of word count for pathology reports. f, The average performance across 7 distinct types of applications, including unimodal (blue boxes) and multimodal tasks (green boxes). g, Critical Difference (CD) Diagram [1] of the average rank performance across 43 diverse tasks for mSTAR and the competing pathology foundation models, in which Bonferroni-Dunn test is employed to assess whether there is a critical difference across different models on multiple datasets. In the CD diagram, a significant difference ($P<0.001$) is indicated by the non-overlapping vertical lines of CD connecting different models.
The recent advancements in foundation models (FMs) [2, 3, 4, 5] for computational pathology (CPath) have demonstrated considerable progress in an incredibly broad spectrum of clinical tasks, such as cancer diagnosis, treatment and prognosis. Despite encouraging performance in general-purpose pathology foundation models, there are still several unresolved challenges.
First, massive multimodal data in line with clinical practices are under-utilized for pretraining, such as pathology reports and gene expression profiles. Existing pathology FMs either focus on vision-only [3] or vision-captions data [2, 4], in which captions are too short to offer sufficiently rich information although attempting to incorporate different modalities. The power of multimodal data has been repeatedly substantiated not only in the general machine learning community [6, 7] but also in the field of medical cancer research [8, 9, 10]. In the clinical workflow, as shown in Figure 1a, pathology reports often provide the complete and exhaustive information of whole slides, while patients’ gene expression profiles offer insights of quantitative molecular dynamics that can complement the qualitative morphological view provided by a slide. The integration of these slide-level multimodal data can establish a broad and holistic perspective, thereby undoubtedly enhancing the capabilities of pathology FMs to perform various clinical tasks.
Second, existing efforts in pathology FMs are primarily directed towards the modelling of patch/ROI-level data [2, 3, 4], leading to limited contexts for slide-level clinical applications. Conventional models typically treat individual patch images as independent samples for pretraining a patch extractor, and subsequently employ multiple instance learning (MIL) [11, 12, 13] to perform slide-level modelling based on embedded patch features. Recent concurrent work has attempted to pretrain a slide-level FM [5]. However, it is achieved by pretraining a slide aggregator on top of pre-extracted patch features with a fixed patch extractor. This way poses an inherent limitation that the upper bound of pretraining performance is inevitably constrained by the quality of patch features, thus limiting their generalization abilities.
In this study, we curated the largest multimodal dataset including three modalities: H&E diagnostic WSIs, pathology reports and RNA-Seq data, resulting in 26,169 slide-level modality pairs (22k used for pretraining) from 10,275 patients across 32 cancer types sourced from The Cancer Genome Atlas (TCGA) (Figure 1c-e). To leverage these multimodal data for CPath, we developed a novel whole-slide pretraining paradigm which injects multimodal knowledge into the pathology FM, termed Multimodal Self-TAught PRetraining (mSTAR) (Figure 1b and  2), broadening the context of modelling into the slide level. To tackle the aforementioned challenges, this paradigm first pretrains a slide aggregator that absorbs multimodal knowledge via slide-level contrastive learning. This slide aggregator will act as a bridge that injects whole-slide contextual multimodal knowledge into the patch extractor at the next stage. Specifically, the slide aggregator gathers all pre-extracted patch features of an entire slide and integrates them into a slide-level representation, which then contrastively aligns with other modalities at the slide level. At the second stage, the pretrained aggregator seamlessly propagates multimodal knowledge learned at the slide-level in the first stage to the patch extractor through self-taught training. During self-taught training, the pretrained aggregator serves as the “Teacher” to supervise the pretraining of the patch extractor, which enforces patch features embedded by the patch extractor to closely resemble the ones re-embedded by the pretrained aggregator.
The proposed paradigm revolutionizes the workflow of pretraining for computational pathology, which allows the pathology FM to possess powerful whole-slide abilities. Furthermore, to the best of our knowledge, this is the first endeavor to inject multimodal knowledge at the slide-level context into pathology foundation models, broadening the scope of contextual modelling from unimodal to multimodal knowledge and from patch-level to slide-level. To systematically investigate the capabilities of mSTAR including unimodal and multimodal applications, we evaluate 7 types of applications on 43 diverse subtasks, resulting in the largest spectrum of downstream tasks. The experimental results have demonstrated that, mSTAR showcased consistent superiority in not only unimodal but also multimodal applications on average, with statistically significant differences compared to previous state-of-the-art (SOTA) FMs (Figure 1f-g). These findings indicate the noteworthy power of multimodal knowledge at the whole-slide context, which endows pathology foundation models with enhanced capabilities that extend beyond the patch-level unimodal context, enabling them to achieve more generalized performance.
Figure 2: The Overview of mSTAR Pipeline. mSTAR is a whole-slide pretraining paradigm comprising two-stage pretraining. a, Stage 1 aims to inject multimodal knowledge into a slide aggregator by slide-level contrastive learning among WSIs, pathology reports and gene expression data. b, Stage 2 aims to seamlessly propagate multimodal knowledge learned at the slide level into the patch extractor by Self-Taught training, which leverages the slide aggregator pretrained in Stage 1 as “Teacher” and enforces patch extractor to be “Student”.
2 Results
2.1 The Overview of mSTAR
The proposed mSTAR aims to provide a novel whole-slide pretraining paradigm that injects multimodal knowledge into the pathology foundation model. Compared with existing pathology foundation models, mSTAR has the following innovative designs to fully unleash its power in a wide spectrum of pathological downstream tasks. First, clinical multimodal data are fully harnessed in pretraining to endow the pathology FM with multimodal knowledge for comprehensive perspectives in clinical tasks. Second, the whole-slide pretraining paradigm provides an alternative way to obtain whole-slide contexts for pathology FMs through self-taught training. To the best of our knowledge, this is the first work to inject multimodal knowledge at the whole-slide context into a pathology FM, broadening the contextual understanding for CPath from patch-level to slide-level and from unimodal to multimodal knowledge. The overview of mSTAR is shown in Figure  2, consisting of two stages of pretraining.
In the first stage, the objective is to inject multimodal knowledge into the slide aggregator by slide-level contrastive learning among three modalities, i.e., WSIs, pathology reports and RNA-Seq profiles. Note that the pretrained slide aggregator will act as a bridge that propagates multimodal knowledge into the patch extractor in the next stage. To this end, as shown in Figure 2a, we first utilized a pretrained patch extractor, a state-of-the-art pathology foundation model named UNI [3], to encode each patch image of a slide into patch features. Then the resulting patch features are fed into a slide aggregator and integrated into a slide-level representation which is subsequently aligned with other modalities through inter-modality contrastive learning. Furthermore, to mitigate the influence of heterogeneity across different types of cancers, the pretraining of the slide aggregator is also supervised by inter-cancer contrastive learning. This approach brings samples of the same cancer type closer together while concurrently pushing samples of different cancer types apart.
In the second stage, the pretrained slide aggregator acquiring multimodal knowledge, can serve as the teacher model to seamlessly propagate multimodal knowledge at the slide-level context into the patch extractor, called Self-Taught Training (Figure  2b). Specifically, the patch extractor is pretrained through encouraging the extracted patch features to be as similar as possible to those re-embedded by the pretrained aggregator. At the same time, to avoid catastrophic forgetting, we also enforce a similarity constraint between the extracted features and ones embedded by the exponential moving average (EMA) patch extractor.
With these two stages, multimodal knowledge at the whole-slide context can be seamlessly embedded into foundation models. As a result, the model acquires the ability to comprehend both patches and the entire WSI, which facilitates downstream tasks at different levels. In the end, the pathology foundation model can achieve advanced abilities with the extended context from patch-level to slide-level and from unimodal to multimodal knowledge. More details of mSTAR can be found in the Section 4.2.
2.2 Pathological Slide Classification for Diagnosis and Treatment
Figure 3: Slide Classification Performance. a, Comparison of Macro-AUC between mSTAR and compared methods on 12 slide classification tasks including 6 diagnostic subtasks and 6 molecular prediction subtasks. They are evaluated with ABMIL and TransMIL trained from scratch, based on different patch extractors. “Overall” refers to the average result between ABMIL and TransMIL. b, The average Macro-AUC on 12 slide classification datasets. c-d, The slide classification performance (Macro-AUC and 95% CI) on 6 diagnostic subtasks and 6 molecular prediction subtasks with ABMIL and TransMIL, respectively. If the performance of mSTAR is the best one compared against the compared FMs, P-value would be presented. mSTAR+ represents mSTAR (patch extractor) equipped with the pretrained aggregator (TransMIL+). The colors of legends are shared across sub-figures a-d. e, The performance (Macro-AUC) gains of the aggregator (TransMIL) between “before” and “after” pretraining among various datasets. “Positive” values indicate performance increases compared with the ones equipped with TransMIL built from scratch, and vice versa. Results of the bar chart in orange and green background are from diagnostic tasks and molecular prediction, respectively. All of performance increases showed significant differences with $P<0.05$.
Slide classification holds significant relevance in clinical settings, particularly in accurate cancer diagnosis and treatment. For example, it aids in categorizing the specific subtype of cancer, which in turn guides the healthcare professionals in precise diagnosis and treatment planning. Additionally, pathologists can identify molecular markers within tissue slides, allowing them to determine the molecular subtypes of cancer. This information is of utmost importance as it forms the foundation for developing personalized treatment plans and enhancing treatment effectiveness. Therefore, the evaluation of slide-level diagnostic tasks, such as cancer subtyping, grading, etc., is the first step of this study, followed by molecular prediction based on the analysis of tissue slides. From the perspective of technique, here we are investigating unimodal abilities through pathological slide classification. To this end, 12 diverse slide classification datasets were used for assessing the ability of decision-making diagnosis at the slide level, consisting of 6 diagnostic subtasks and 6 molecular prediction subtasks. For diagnostic tasks, the following subtasks were included: breast metastasis detection on CAMELYON [14, 15], prostate ISUP grading on PANDA [16], ovarian cancer subtyping on UBC-OCEAN [17, 18], BRCA subtyping on TCGA BRCA [19], NSCLC subtyping on TCGA NSCLC [19] and RCC subtyping on RCC-DHMC [20]. For molecular prediction, we performed the prediction of molecular subtypes ER, HER2 and PR on BCNB [21] datasets, respectively. This task also includes molecular subtypes prediction on TCGA-BRCA, TCGA-CRC and TCGA-GBMLGG datasets [19]. The details of every task are illustrated in the Section 4.3.
The evaluations are two-fold: 1) we examined patch features of different pathology foundation models in slide classification (Figure 3a-d), including ResNet50 (R50) [22], PLIP [2], CONCH [4], UNI [3] and our mSTAR. To obtain slide/patient-level classification decision, we adopted two commonly used MIL backbones as the slide aggregator trained from scratch, attention-based MIL (ABMIL) [11] and TransMIL [13] for integrating these extracted patch features. ABMIL is a simple yet robust MIL approach, which is usually used for evaluation in previous foundation model research [3, 4, 5]. Considering that we employed the pretrained TransMIL as the “Teacher” in the proposed method, we also present the performance of TransMIL from scratch for all extractors. 2) Since we pretrained a TransMIL in the first stage in the proposed paradigm, we additionally explore how it collaborates with the pretrained patch extractor (Figure 3d), leading to an advanced version (mSTAR+). We further compared the performance of different extractors between being equipped with a non-pretrained TransMIL (built from scratch) and the pretrained TransMIL+ (Figure 3e).
All comparisons are based on the metric of Macro-AUC, a commonly used classification metric, which does not rely on the selection of the decision threshold and is insensitive to the sample ratio of various classes. Furthermore, to avoid the impact of potential inherent bias in various datasets and different MIL backbones, following the previous research [1], the average rank of Macro-AUC was reported as well (Figure 3a) and the one-sided Wilcoxon signed-rank test was performed on various datasets to examine the statistical difference between mSTAR and other FMs.
From an overall perspective, we assessed the average rank and average Macro-AUC for mSTAR and the compared foundation models across 12 diverse subtasks. The results showed that mSTAR consistently achieved the first place across different tasks, regardless of different MIL backbones employed or when considering the overall performance. For the consistency of performance improvements, Figure 3c-d (Extended Data Table 2-3) demonstrated that mSTAR achieved the best performance on 11 out of 12 subtasks with ABMIL and 9 out of 12 subtasks with TransMIL. When we incorporated the pretrained aggregator, TransMIL+, mSTAR+ won other SOTA FMs on 10 out of 12 subtasks (Extended Data Table 4).
When we delve into the specific metric of Macro-AUC, we found that mSTAR achieved the best performance with increases of +1.3% overall ($P<0.001$) compared to the second-best model, UNI, as shown in Figure 3b. The upward trends of average Macro-AUC were witnessed in ABMIL and TransMIL, going from 0.905 to 0.908 ($P<0.001$) and from 0.870 to 0.892 (+2.2%, $P<0.005$), respectively. These results indicate that mSTAR has the promising ability of slide classification in all. More detailed results for each subtask are presented as follows.
mSTAR improves Slide-level Diagnostic Tasks. Among diagnostic tasks, the consistent improvements in ABMIL become apparent, as mSTAR outperformed the second-best approach on all diagnostic datasets with significant differentiation ($P<0.001$, Figure 3c). TransMIL exhibited a similar trend (Figure 3d), in which results on 4 out of 6 datasets indicated significant differences.
When considering the incorporation with the pretrained aggregator, mSTAR+ exhibited further advancements on various datasets and achieved the best on 5 out of 6 diagnostic tasks. For example, compared to the second-best compared FM, mSTAR+ outperformed by +2.01% on CAMELYON ($P<0.001$). Furthermore, an interesting finding from Figure 3e (Extended Data Table 4) was that the pretrained aggregator was even able to collaborate with other patch extractors achieving critical performance gains, although it was solely pretrained on features extracted by mSTAR. As shown in Figure 3e, it consistently displayed superior performance in the majority of tasks for different FMs. Among these, the pretrained aggregator achieved a notable increase of up to +9.10% ($P<0.001$) compared to the one built from scratch, although there were slight decreases (-1.15% at most) of performance in some specific subtasks. From Figure 3e, we also observed that the enhancements made to CONCH, UNI, and mSTAR were not substantial, which could be attributed to the fact that their performances have already approached 100% and the room for improvement was significantly limited on these 6 subtasks.
mSTAR improves Molecular Prediction. Following the same evaluation strategy in diagnostic tasks, we also investigated its capability of molecular prediction from tissue slides.
From the perspective of consistency, mSTAR with ABMIL achieved the best performance on 5 out of 6 tasks (Figure 3c and Extended Data Table 2), in which 4 tasks exhibited a significance difference ($P<0.001$). With TransMIL, clearer improvements (Figure 3d and Extended Data Table 3) were observed that 4 out of 6 tasks demonstrated a significant difference ($P<0.05$). Taking a closer look at Macro-AUC, mSTAR outperformed other SOTA foundation models by +2.23% overall with a critical difference ($P<0.001$) on average across different MIL backbones (Extended Data Table 2-3). Specifically, for TransMIL, a substantial increase of +3.71% on average Macro-AUC across 6 molecular tasks was observed, although mSTAR achieved slight improvement with ABMIL ($P<0.05$). In particular, substantial performance increases of +4.64% and +4.21% were observed on BRCA-Molecular and CRC-Molecular tasks ($P<0.001$), respectively.
In addition, when incorporating the pretrained aggregator TransMIL+ (Extended Data Table 4), further improvements were achieved where mSTAR+ outperformed the compared FMs on 5 out of 6 datasets (Figure 3d, $P<0.001$) by +6.35% on CRC-Molecular, +2.93% on BCNB-HER2, +1.82% on GBMLGG-IDH1 and +1.5% on BCNB-ER. From Figure 3e, we observed a similar trend that the majority of subtasks demonstrated further enhancements for different extractors. An interesting finding different from that of diagnostic tasks was that performance gains on CONCH and UNI, in which there were no notable increases in diagnostic tasks, demonstrated remarkable improvements in molecular prediction (e.g., +13.10% on CRC-Molecular and +7.70% on BRCA-Molecular for CONCH, +9.00% on BRCA-Molecular and +6.29% on BCNB-ER for UNI, $P<0.001$). This validates that the involvement of RNA-Seq data in pretraining contributes to facilitating molecular prediction, as pretraining data used in CONCH and UNI did not include gene expression profiles and our aggregator was contrastively pretrained on RNA-Seq data. It is worth noting that the pretrained aggregator provided relatively modest gains for our extractor mSTAR (+2.66% on GBMLGG-IDH1, +2.14% on CRC-Molecular and +1.73% on BCNB-HER2, $P<0.001$). This can be explained by the fact that the knowledge regarding gene expression profiles has already been incorporated into the patch extractor mSTAR through self-taught training, which has resulted in satisfactory performance in molecular prediction.
Overall, mSTAR demonstrates consistent performance gains on both diagnostic tasks and molecular prediction at the slide-level with a significant difference, compared with prior patch-level SOTA pathology foundation models. This suggests that the modelling of whole-slide contexts is conducive to slide-level classifications. Furthermore, incorporating the pretrained aggregator is able to result in further improvements, especially on molecular prediction tasks. In particular, when using the aggregator pretrained on three modalities, those extractors that have not been exposed to gene expression profiles can acquire additional significant enhancements in molecular prediction. This can be attributed to the incorporation of RNA-Seq data during the pretraining process, thereby providing additional enhancements.
2.3 Pathological Survival Analysis for Prognosis
Figure 4: Pathological Survival Analysis Performance. a, Comparison of C-Index between mSTAR and compared methods on 9 TCGA datasets. They are evaluated with ABMIL and TransMIL trained from scratch, based on different patch extractors. “Overall” refers to the average results between ABMIL and TransMIL. b, The average C-Index on 9 TCGA datasets. c-d The survival analysis performance (C-Index and 95% CI) on 9 TCGA datasets with ABMIL and TransMIL, respectively. If the performance of mSTAR is the best one compared against the compared FMs, P-value would be presented. mSTAR+ represents mSTAR (patch extractor) equipped with the pretrained aggregator (TransMIL+). The colors of legends are shared across sub-figures a-d. e The performance (C-Index) gains of the aggregator (TransMIL) between “before” and “after” pretraining among 9 TCGA cancer types. “Positive” values indicate performance increases compared with the ones equipped with TransMIL built from scratch, and vice versa. All of performance increases showed significant differences with $P<0.05$, except the ones marked by $\times$.
Prognostic analysis is an intricate clinical endeavor, necessitating a thorough analysis from a multitude of facets. In this regard, the integration of multimodal data has proven instrumental in enabling more comprehensive prognostic assessments [8, 9, 10, 23]. Therefore, it is crucial to delve into the exploration of the role of multimodal knowledge within the broader whole-slide context in enhancing prognosis estimation. In this study, we assessed a key prognostic task, overall survival (OS) prediction, on top of pathological tissue slides.
mSTAR advances Cancer Survival Prediction. We investigate the capability of predicting overall survival (OS) across 9 diverse cancers, based on the metric of C-Index, a metric commonly used in survival prediction. To avoid data leakage, the data used for evaluation were held out from pretraining data. To ensure the reliable evaluation, we excluded cancer types with fewer than 400 cases or those with too few uncensored data out of the 32 cancer datasets in TCGA [19], resulting in 9 cancer datasets: Breast Invasive Carcinoma (BRCA), Colon Adenocarcinoma and Rectum Adenocarcinoma (CRC), Glioblastoma Multiforme and Brain Lower Grade Glioma (GBMLGG), Head and Neck Squamous Cell Carcinoma (HNSC), Kidney Renal Clear Cell Carcinoma (KIRC), Lung Adenocarcinoma (LUAD), Lung Squamous Cell Carcinoma (LUSC), Skin Cutaneous Melanoma (SKCM) and Uterine Corpus Endometrial Carcinoma (UCEC). Similarly, following the settings of slide-level classification, we also used ABMIL and TransMIL for evaluating patch extractors and then the performance of the pretrained slide aggregator was assessed. More details are presented in Section 4.3.
For consistency of performance increases, mSTAR performed best compared to other foundation models, where Figure 4a shows that mSTAR always ranked in front of other foundation models for different MIL backbones and the overall performance. When considering each individual subtask, mSTAR achieved the best performance on 7 out of 9 datasets with ABMIL (Figure 4c and Extended Data Table 5), in which 5 tasks demonstrated a significant difference ($P<0.01$) compared to the second-best approach. With TransMIL (Figure 4d and Extended Data Table 6), mSTAR presented more consistent improvements, which outperformed compared approaches on 8 out of 9 cancer types and showed significant improvements compared to the second-best method in 7 cancer types ($P<0.001$).
When specifically assessing the C-Index metric, the average C-Index showcased significant performance gains compared against the second-best model (UNI) by +1.98% for TransMIL ($P<0.005$), +0.53% for ABMIL ($P<0.05$), and +1.30% for the average performance of different MIL backbones (Overall, $P<0.001$). Next, we delved into details of performances across 9 datasets, where we observed apparent improvements ($P<0.001$) of +3.7% on SKCM, +3.34% on LUAD, +2.87% on LUSC and +1.03% on UCEC over the second-best model based on TransMIL (4d and Extended Data Table 6). These results indicated the effectiveness of multimodal knowledge involved in mSTAR in extracting discriminative features for distinguishing complex survival patterns in pathological images.
We also investigated the performance of incorporating the pretrained aggregator. From observations of Figure 4d, mSTAR+ equipped with the pretrained aggregator pushed performance to higher levels (+5.23% on SKCM, +4.9% on UCEC, +2.96% on BRCA, +2.60% on KIRC and +2.33% on CRC, $P<0.001$) compared to the second-performing approaches. By comparing TransMIL “before” and “after” pretraining for various extractors (Figure 4e, Extended Data Table 7), we observed that, in the majority of cases, the pretrained aggregator exhibited superior performance, surpassing the one built from scratch by a margin of up to +7.79% ($P<0.001$). This indicates the introduction of pathology report and gene expression profiles is beneficial to survival analysis by providing additional expert knowledge and molecular information. In addition, a similar trend to that of slide classification was observed that, PLIP and CONCH that had a lack of knowledge of gene expression profiles exhibited more notable improvements than mSTAR, when collaborating with our aggregator pretrained on gene data. This supports the analysis illustrated in the results of slide classification again, in which multimodal knowledge has been transferred into the patch extractor mSTAR, leading to relatively moderate improvements when using the pretrained aggregator. In other words, mSTAR itself has absorbed the knowledge in pathology reports and gene expression profiles, which has contributed to survival prediction.
To sum up, mSTAR demonstrated similar trends on both slide classification and survival analysis with consistent improvements, which repeatedly validated its capability on unimodal pathological clinical tasks. Additional results on the pretrained aggregator, on the one hand, provided extra evidence of the significance of pathology reports and RNA-Seq data at the whole-slide context, especially on molecular prediction and survival analysis. On the other hand, it indicated that mSTAR itself could acquire multimodal knowledge at the slide level, thereby resulting in better performance in these diverse tasks.
2.4 Multimodal Capability
Figure 5: Multimodal Survival Analysis Performance. a, Comparison of C-Index between mSTAR and compared methods on 9 TCGA datasets. They are evaluated with different multimodal fusion models, MCAT, Porpoise, MOTCat and CMTA trained from scratch based on the compared patch extractors and mSTAR. “Overall” refers to the average results among these multimodal fusion methods. b, The average C-Index on 9 TCGA datasets. c-f The survival analysis performance (C-Index and 95% CI) on 9 TCGA datasets with various multimodal fusion models, respectively. If the performance of mSTAR is the best one compared against the compared FMs, P-value would be presented. The colors of legends are shared across all sub-figures.
Figure 6: Performance of Few-shot Slide Classification. a, Few-shot Slide Classification via Prototypical MI-FewShot. Top-K Patches of slides in the support set are selected to construct class prototypes for each category. They are chosen by higher similarities to the class embedding of class prompts via the pretrained text encoder. For a query slide, similarities between all patches and each class prototype are computed, and the class label is assigned to the most similar class prototype by top-K patches voting. b, Average Rank on Macro-AUC of compared FMs and mSTAR on 6 subtasks and overall. “Overall” refers to the average result among various subtasks. c, Average Macro-AUC across k-shots on 6 subtasks and overall. The numbers in green indicate the performance improvement over the vision-only FM, UNI. d-i, Macro-AUC across varying shots on every dataset for each FM. The detailed performance is presented for mSTAR and the best or second-best FM. The colors of legends are shared across sub-figures b-i.
Multimodal contrastive pretraining naturally allows the model to be applied to multimodal downstream tasks. In this study, since mSTAR aligned multimodal data that may contribute to alleviating inter-modal heterogeneity, we first examine whether features extracted by mSTAR facilitate multimodal fusion by assessing multimodal survival analysis tasks. Next, we investigate more cross-modal abilities, including slide-level few-shot and zero-shot classification and pathological report generation.
mSTAR facilitates Multimodal Fusion. The involvement of multimodal data in the pretraining process can enhance the model’s capability of capturing complex interactions across different modalities and aligning each other by contrastive learning, thereby potentially alleviating the heterogeneity of different modalities and improving its performance on a wider range of cancer types. To validate this, we replaced pathological features with ones extracted by various extractors in existing multimodal fusion models for 9 TCGA cancer survival prediction tasks, to observe the differences that would arise. Specifically, recent SOTA multimodal fusion models were employed in this study, including MCAT [8], Porpoise [24], MOTCat [9] and CMTA [10].
On the whole, mSTAR has clearly outperformed other SOTA methods by a wide margin. Considering average rank, mSTAR ranked between 1.22 and 1.67 among various fusion models and the overall rank is 1.47, which left the second-best approach UNI far behind (Figure 5a) ranking at 2.56 to 3.00 and 2.68 on average. For average C-Index (Figure 5b), mSTAR achieved consistent and notable enhancement in multimodal fusion with a significant difference, with average performance increases of +1.8% ($P<0.001$). Among different multimodal fusion models, mSTAR outperformed the second-best UNI by +1.7% ($P=0.02$) for MCAT, +1.1% ($P=0.01$) for Porpoise, +2.5% ($P<0.005$) for MOTCat and +1.9% ($P<0.005$) for CMTA.
Across various datasets, based on MCAT, mSTAR surpassed other SOTA approaches on 7 out of 9 tasks (Figure 5c and Extended Data Table 8), especially on LUAD (+3.3%, $P<0.001$) and BRCA (+1.9%, $P<0.001$). In the case of MOTCat, mSTAR excelled in 7 out of 9 tasks (Figure 5d and Extended Data Table 10) with performance increases of 0.6%-4.8% ($P<0.001$). mSTAR with Porpoise demonstrated superior performance in the majority of tasks, topping 5 out of 9 datasets (Figure 5e and Extended Data Table 9), which increased the second-best model by up to +3.8% ($P<0.001$). For CMTA, mSTAR delivered the highest performance in 7 of 9 (Figure 5f and Extended Data Table 11), advancing the second-best one by 0.4%-4.1% ($P<0.001$).
In a nutshell, the remarkable increases across various datasets and diverse multimodal fusion backbone models vividly demonstrate the tremendous contributions of multimodal knowledge embedded by slide-level multimodal contrastive learning in facilitating multimodal fusion.
mSTAR advances Few-shot Slide Classification.
Figure 7: Performance of Zero-shot Slide Classification. a, Zero-shot Slide Classification via MI-Zero [25]. For a query slide, similarities between all patches and each class embedding are computed, and the class label is assigned to the most similar class prototype by top-K patches voting. b, Average Rank on Macro-AUC of compared FMs and mSTAR across 6 subtasks. “Overall” refers to the average result among various subtasks. c, Macro-AUC on 6 subtasks and overall. The numbers in green indicate the improvement over the vision-only FM, UNI. The colors of legends are shared across sub-figures b-c. If the performance of mSTAR is the best one compared against the compared FMs, P-value would be presented.
Few-shot learning aims to achieve accurate prediction using only a small number of labeled data for each class, which is especially valuable in the context of medical scenarios where data scarcity is a common challenge. The ability of generalizing to new tasks with a few labels can be enhanced by multimodal data by offering a more comprehensive and diverse set of information. In this study, we investigated the few-shot capability of different patch extractors given different shots of labeled slides.
To ensure that the pretrained extractor can solely rely on the limited labeled slides provided, we excluded TCGA classification subtasks for evaluation, since mSTAR was pretrained on TCGA data. As a result, every patch extractor was assessed on 6 slide classification subtasks, CAMELYON, PANDA, UBC-OCEAN, BCNB-ER, BCNB-PR and BCNB-HER2. To enable the few-shot classification in the context of multiple instance learning, we proposed MI-Fewshot, as illustrated in Figure 6a and the Section 4.3. It utilizes a pretrained text encoder to select the most relevant patches from a limited amount of labeled data using similarity scores. These selected patches are then averaged to form the class prototype for accomplishing few-shot classification. Since MI-Fewshot requires a pretrained text encoder, we equipped those vision-only extractors, i.e., R50 and UNI, with the pretrained text encoder of mSTAR for assessment, while other extractors used their own designated text encoders.
Across 6 subtasks, mSTAR ranked the 1st place overall and achieved the best performance on 4 subtasks (Figure 6b and Extended Data Table 12), where mSTAR outperformed second-best FMs on average by a large margin (+4.1% on Camelyon, +1.0% on UBC-OCEAN, +2.0% on BCNB-PR and +0.9% for overall), as shown in Figure 6c (Extended Data Table 12). To demonstrate the benefits of multimodal knowledge, we also presented the performance increase compared to the vision-only SOTA FM, UNI. On 4 out of 6 subtasks, mSTAR exhibited better performance on average, with an additional increase on PANDA by +1.7%.
Following previous studies [4], we examined performance on varying shots of every subtask (Figure 6d-i and Extended Data Table 13-18). We observed that mSTAR has showcased superior performance with a larger margin when given fewer slides on Camelyon, while the performance gap has gradually diminished as the number of slides increased. This indicated mSTAR performed better on the setting with fewer slides. On UBC-OCEAN, mSTAR always performed the best. For PANDA, mSTAR consistently outperformed vision-only UNI across varying shots, which also indicated the effectiveness of multimodal knowledge.
All in all, these results demonstrated mSTAR advanced generalization capabilities in label-efficient scenarios, which could be attributed to slide-level multimodal contrastive learning. Specially, by comparing with vision-only FMs, mSTAR showcased clear superiority, which indicated the extra benefit of multimodal knowledge.
Figure 8: Pathology Report Generation Performance. a, quantitative results (BLEU-1-4, METEOR and ROUGE-L) of various compared models and mSTAR. b, an example of report generation, where texts in red boxes are grount-truth from doctors and ones in dark green boxes are generated with pathological features from various models. The words highlighted in green are aligned with ground-truth, while ones marked in red are missing by mSTAR from ground-truth.
mSTAR enables Zero-shot Slide Classification. In the context of open clinical practice, out-of-the-box capability of diagnostic decision-making without the need for task-specific fine-tuning is crucial for efficient utilization of multimodal models, especially in resource-constrained scenarios where the access to sufficient data and computational resources may be limited, such as underdeveloped areas. Therefore, we also evaluated zero-shot’s capability of each extractor on 6 slide classification subtasks where we excluded TCGA subtasks as well, following the similar settings in few-shot’s evaluation. To produce slide-level predictions for patch extractors, MI-Zero [25] was adopted through top-K patches voting based on patch similarities to class prototypes embedded by the pretrained text encoders (Figure 7a).
Across 6 slide subtasks, mSTAR outperformed other FMs on half of the tasks and performed best on the overall result (Figure 7b and c). Similarly, we also compared mSTAR with the vision-only FM, UNI, to validate the effectiveness of multimodal knowledge injection. As a result, notable performance increases were observed on 4 of 6 tasks, and meanwhile mSTAR achieved a comparable performance with UNI on an extra UBC-OCEAN task. Specifically, compared to the second-best FM, UNI, mSTAR achieved clear enhancement in these tasks by +3.1% on average ($P<0.001$) with a significant difference. In particular, on BCNB-ER task, a remarkable rise of +9.4% ($P<0.001$) was observed compared to the second-best FM, UNI, and we further saw performance enhancement over UNI from 0.618 to 0.725 (+10.7%, $P<0.001$). For CAMELYON and BCNB-PR, there were notable boosts of +5.7% and +4.0% over the second-performing FM ($P<0.001$), respectively. For PANDA, mSTAR significantly enhanced UNI from 0.45 to 0.462 ($P<0.001$).
Overall, despite not being the best performing on all subtasks, mSTAR clearly outperformed the second-best FM with a large margin on half of the subtasks. When compared with the vision-only FM, mSTAR performed better or comparative on 5 out of 6 subtasks, which indicates the effectiveness of multimodal contrastive learning. These results showcased that mSTAR can achieve notable advancements in its out-of-the-box ability on certain subtasks. However, there is still considerable room for further enhancements. The integration of additional multimodal data through the pretraining approach of mSTAR has the potential to enhance its zero-shot capability.
mSTAR excels in Pathological Report Generation. Automated generation of pathology reports has enormous potential in simplifying the report-writing process and reducing the workload burden on pathologists. Next, we assessed its capability of report generation on TCGA datasets. It is worth mentioning that cases used for evaluation are completely held out from the pretraining materials. We presented both quantitative and qualitative results.
From the quantitative perspective, we evaluated multiple metrics including BLEU, METEOR and ROUGE-L to assess various aspects of the generated text, such as precision of n-grams (contiguous sequences of words), order, alignment, recall, etc. Across different metrics, mSTAR consistently outperformed the second-best approach by nearly 1% (Figure 8a and Extended Data Table 20), which indicates better quality of generated reports.
We continued to qualitatively evaluate the quality of generated reports. A case study is presented in Figure 8b and more cases are provided in the Extended Data Figure 10. From the case of Figure 8b, we found that mSTAR is able to more accurately generate important diagnostic facts with minimal omissions, such as clinical diagnosis. The result of Figure 8b generated by CONCH, PLIP and R50 contain a significant number of factual errors. Compared with the second-best approach, UNI, mSTAR is capable of generating immunohistochemical results on top of H&E WSIs, such as HER2, which may be attributed to the involvement of gene expression data in pretraining. This may align with the earlier investigation of mSTAR’s superior performance in molecular prediction. From the case of Extended Data Figure 10a, mSTAR is able to generate the majority of key diagnostic information, while other FM just produced a few key words. The case of Extended Data Figure 10b demonstrated that mSTAR excelled in producing precise histologic type and grade, while the majority of the results generated by other FMs are factually incorrect. Although UNI included grading information in the case of Extended Data Figure 10c, the result of pathological T staging was significantly erroneous. This can be attributed to the involvement of gene expression profiles, since RNA-Seq data are known to be strong indicators of staging [26, 27], subtyping [28] and grading [29, 30]. However, all FMs are unable to provide accurate quantity estimations, such as lesion sizes and tumor location. The possible reasons for the failure of quantity estimations are that they lack information regarding the magnification and locations of target objects mentioned in a certain report. These pieces of information are likely recorded by doctors during the specimen collection process and cannot be directly inferred from the WSI.
In conclusion, the multimodal data endows mSTAR with unique knowledge from each modality, such as identifying molecular profiles, enabling it to generate more comprehensive and accurate reports. These encouraging results underscored the potential of mSTAR for report generation and its significance in pathological decision-making.
3 Discussion
In this study, we delve into how to harness the full potential of multimodal data to effectively advance the performance of the pathology foundation models and bridge the gaps of previous foundation models in multimodal capabilities. Additionally, we explored a new whole-slide pretraining paradigm for CPath, which broadened the context of modelling for better performance on slide-level tasks. We curated the largest multimodal dataset consisting of WSIs, pathology reports and RNA-Seq data, which contained over 26k public modality pairs from 10,275 patients across 32 major cancer types. Diverse experimental results on 43 subtasks demonstrated that mSTAR excelled in not only unimodal tasks but also multimodal tasks at the slide level. In particular, with the inclusion of gene expression data, mSTAR performed well in slide-level molecular prediction. Furthermore, multimodal contrastive learning facilitated multimodal fusion tasks and endowed the model with more generalized capabilities in low-shot scenarios. Simultaneously, the incorporation of clinical pathology reports in the pretraining process resulted in the advanced capability of report generation by providing the expert knowledge.
In the realm of prior investigations into pathology foundation models, two prominent categories have emerged: vision-only models [3, 31, 5] and vision-language models [4, 2]. However, these approaches fail to tap into a vast wealth of information inherent in clinical pathology reports and gene expression profiles. Pathology reports usually provide authentic expert knowledge in line with the clinical practice, while gene expression profiles always serve as robust indicators of disease status for clinical applications in diagnosis [32] and prognosis [33]. As shown in Extended Data Table 1, the involvement of pathology reports and gene expression data can bring extra performance gains. We also found that there was still considerable room for further improvement in molecular prediction, multimodal fusion for survival analysis and report generation in existing FMs, which can benefit from the inherent knowledge in pathology reports [34] and gene expression profiles [29, 28, 30]. Simultaneously, we notice that the majority of existing FMs primarily focus on patch/ROI-level models and short texts, in which restricted contexts hinder their practical performance in slide-level clinical applications, such as survival analysis and report generation. It is worth noting that, unlike CONCH supervised by generative loss, mSTAR without involving generative components in pretraining, still demonstrated encouraging performance in report generation with producing more comprehensive texts. Recently, beyond working on small patches/ROIs, we noticed that a study attempted to work on slide-level foundation models, which pretrained the model on patch features. However, the pretrained performance significantly depends on the quality of patch features. In other words, their performance would be limited by the patch extractor. We believe that end-to-end pretraining is a promising solution in the future, while its current implementation is hindered by hardware limitations. Therefore, mSTAR bridges this gap through self-taught training to seamlessly transfer the knowledge captured by the slide aggregator into the patch extractor.
Distinct from previous researches, our study provides the uniqueness in two folds. First, our findings showcase the remarkable power of leveraging multimodal data to drive the progress of pathology foundation models, especially in enhancing multimodal capabilities. Second, we found a unified way to bridge the gap between slide-level and patch-level pretraining, bringing us closer to achieving end-to-end pretraining on raw slide data. We believe this innovative unified paradigm will revolutionize the workflow of pretraining for CPath. Moreover, this paradigm allows the injection of multimodal knowledge into pathology foundation models in an appropriate manner, which may hold the potential to harness more modalities to construct a stronger foundation model for CPath.
Although preliminary results are encouraging, this study still has several limitations. First, a key limitation is the scale of data for pretraining compared to previous works of pathology foundation models. By expanding the scale of multimodal data for pretraining, we can expect to unlock further potential for enhancing various abilities, such as multimodal capabilities. Second, we still potentially have a long way to go before achieving the true end-to-end foundation model. Before that, mSTAR will serve as an alternative solution to seamlessly bridge slide-level and patch-level pretraining. However, there are still several challenges to be further explored, such as the appropriate way to propagate the pretrained knowledge embedded in the slide aggregator and the architectural design of slide aggregator. In mSTAR, due to a large number of patches of a WSI that would lead to extremely high computational costs, we selected TransMIL with linear time complexity as the slide aggregator. However, the increase in training speed comes at the expense of sacrificing a portion of the performance. Fortunately, a multitude of innovative architectures for handling long sequences are emerging, such as Mamba [35], LongNet [36], etc, which we explore in concurrent work [37]. We believe that these new architectures will undoubtedly create new avenues in exploring more efficient and powerful pretraining paradigms for CPath. In the future, we plan to incorporate more multimodal data into pretraining, such as multi-omics data, and explore new efficient pretraining architecture.
Appendix A Extended Data
(a) TCGA-D1-A179
(b) TCGA-BS-A0V6
(a) TCGA-DF-A2L0
Figure 10: Pathology Reports Generation for several cases.
Table 1: Comparison of the aggregator pretrained on different combinational modalities (Pathology, Reports and RNASeq) for survival analysis. Note that pathological patch features are extracted by UNI. Average C-Index and its std across 9 TCGA survival datasets are reported.
Aggregator (TransMIL) Pretrained Data Avg C-Index
Pathology Reports RNASeq
w/o pretraining 0.666±0.068
w/ pretraining 0.679±0.057
0.679±0.059
0.689±0.055
Table 2: Macro-AUC of Slide Classification with ABMIL on 12 tasks. Best performing model for each metric is bolded. 95% CI is included in parentheses.
Task R50 PLIP CONCH UNI mSTAR(Ours)
Diagnosis CAMELYON 0.9236 (0.8905,0.9567) 0.9242 (0.8915,0.9569) 0.9823 (0.9654,0.9992) 0.9814 (0.9683,0.9945) 0.9934 (0.9865,1.0000)
PANDA 0.8872 (0.8811,0.8933) 0.9045 (0.8988,0.9102) 0.9220 (0.9169,0.9271) 0.9455 (0.9410,0.9500) 0.9468 (0.9423,0.9513)
UBC-OCEAN 0.8940 (0.8644,0.9236) 0.9453 (0.9230,0.9676) 0.9724 (0.9606,0.9842) 0.9732 (0.9612,0.9852) 0.9764 (0.9637,0.9891)
TCGA-BRCA 0.7560 (0.6876,0.8244) 0.8533 (0.8047,0.9019) 0.9386 (0.9119,0.9653) 0.9391 (0.9038,0.9744) 0.9441 (0.9110,0.9772)
TCGA-NSCLC 0.9111 (0.8876,0.9346) 0.9395 (0.9193,0.9597) 0.9723 (0.9613,0.9833) 0.9695 (0.9558,0.9832) 0.9730 (0.9607,0.9853)
RCC-DHMC 0.9589 (0.9440,0.9738) 0.9889 (0.9820,0.9958) 0.9835 (0.9727,0.9943) 0.9889 (0.9818,0.9960) 0.9898 (0.9837,0.9959)
Molecular Prediction BCNB-ER 0.7974 (0.7466,0.8482) 0.8460 (0.8058,0.8862) 0.8503 (0.8072,0.8934) 0.8886 (0.8506,0.9266) 0.8887 (0.8519,0.9255)
BCNB-HER2 0.6683 (0.6075,0.7291) 0.6789 (0.6203,0.7375) 0.7465 (0.6938,0.7992) 0.7409 (0.6890,0.7928) 0.7419 (0.6896,0.7942)
BCNB-PR 0.6889 (0.6289,0.7489) 0.8041 (0.7520,0.8562) 0.8092 (0.7592,0.8592) 0.8297 (0.7787,0.8807) 0.8335 (0.7829,0.8841)
BRCA-Molecular 0.7017 (0.6637,0.7397) 0.7353 (0.6924,0.7782) 0.7506 (0.7161,0.7851) 0.7613 (0.7235,0.7991) 0.7646 (0.7279,0.8013)
CRC-Molecular 0.6304 (0.5830,0.6778) 0.7451 (0.7016,0.7886) 0.8382 (0.7998,0.8766) 0.8548 (0.8227,0.8869) 0.8560 (0.8227,0.8893)
GBMLGG-IDH1 0.8787 (0.8250,0.9324) 0.9345 (0.9022,0.9668) 0.9777 (0.9644,0.9910) 0.9891 (0.9820,0.9962) 0.9902 (0.9831,0.9973)
Avg Macro-AUC 0.8080 0.8583 0.8953 0.9052 0.9082
Avg Rank 5.00 3.88 2.75 2.29 1.08
Table 3: Macro-AUC of Slide Classification with TransMIL on 12 tasks. Best performing model for each metric is bolded. 95% CI is included in parentheses.
Task R50 PLIP CONCH UNI mSTAR(Ours)
Diagnosis CAMELYON 0.9164 (0.8819,0.9509) 0.9161 (0.8808,0.9514) 0.9577 (0.9308,0.9846) 0.9719 (0.9550,0.9888) 0.9689 (0.9446,0.9932)
PANDA 0.8870 (0.8809,0.8931) 0.8929 (0.8868,0.8990) 0.9108 (0.9053,0.9163) 0.9311 (0.9264,0.9358) 0.9317 (0.9270,0.9364)
UBC-OCEAN 0.8891 (0.8536,0.9246) 0.9106 (0.8836,0.9376) 0.9589 (0.9409,0.9769) 0.9643 (0.9482,0.9804) 0.9722 (0.9595,0.9849)
TCGA-BRCA 0.7553 (0.6936,0.8170) 0.8511 (0.8005,0.9017) 0.9335 (0.9061,0.9609) 0.9453 (0.9122,0.9784) 0.9480 (0.9141,0.9819)
TCGA-NSCLC 0.9035 (0.8786,0.9284) 0.9321 (0.9111,0.9531) 0.9699 (0.9564,0.9834) 0.9725 (0.9609,0.9841) 0.9647 (0.9473,0.9821)
RCC-DHMC 0.9674 (0.9541,0.9807) 0.9622 (0.9451,0.9793) 0.9850 (0.9768,0.9932) 0.9902 (0.9843,0.9961) 0.9918 (0.9869,0.9967)
Molecular Prediction BCNB-ER 0.7508 (0.6953,0.8063) 0.8124 (0.7618,0.8630) 0.8705 (0.8297,0.9113) 0.7993 (0.7478,0.8508) 0.8791 (0.8379,0.9203)
BCNB-HER2 0.4955 (0.4355,0.5555) 0.6638 (0.6028,0.7248) 0.7035 (0.6447,0.7623) 0.6916 (0.6338,0.7494) 0.7155 (0.6606,0.7704)
BCNB-PR 0.6327 (0.5670,0.6984) 0.7178 (0.6586,0.7770) 0.8062 (0.7564,0.8560) 0.7832 (0.7293,0.8371) 0.8029 (0.7537,0.8521)
BRCA-Molecular 0.6476 (0.6076,0.6876) 0.7174 (0.6749,0.7599) 0.6773 (0.6338,0.7208) 0.6638 (0.6222,0.7054) 0.7638 (0.7299,0.7977)
CRC-Molecular 0.6214 (0.5675,0.6753) 0.6147 (0.5633,0.6661) 0.6914 (0.6422,0.7406) 0.7765 (0.7349,0.8181) 0.8186 (0.7784,0.8588)
GBMLGG-IDH1 0.9259 (0.8916,0.9602) 0.9282 (0.8953,0.9611) 0.9639 (0.9463,0.9815) 0.9640 (0.9419,0.9861) 0.9556 (0.9276,0.9836)
Avg Macro-AUC 0.7807 0.8275 0.8682 0.8702 0.8924
Avg Rank 4.83 3.92 2.50 2.25 1.50
Table 4: Macro-AUC of Slide Classification with the pretrained TransMIL on 12 datasets. The performance gains of the aggregator (TransMIL) between “before” and “after” pretraining. “Positive”(+) values indicate performance increases ($\uparrow$) compared with the ones equipped with TransMIL built from scratch, and vice versa. The best improvement is bolded.
Task R50 PLIP CONCH UNI mSTAR (Ours)
Diagnosis CAMELYON -0.16% -1.15% -0.50% +1.60% +2.31%
PANDA +0.21% -0.61% -0.03% +0.95% +0.86%
UBC-OCEAN +2.03% +3.19% +0.24% +0.55% -0.41%
TCGA-BRCA +9.10% -0.84% -0.31% -0.91% -0.52%
TCGA-NSCLC +2.07% +1.18% -1.12% +0.26% +1.06%
RCC-DHMC +3.64% +1.54% +1.54% +1.01% -0.05%
Molecular Prediction BCNB-ER +7.27% -0.56% -2.19% +6.29% +0.64%
BCNB-HER2 +12.28% +1.76% +3.39% +5.08% +1.73%
BCNB-PR +9.86% +0.64% -1.70% +1.40% -0.92%
BRCA-Molecular +3.06% -0.46% +7.70% +9.00% -0.83%
CRC-Molecular -0.67% +14.94% +13.10% +4.85% +2.14%
GBMLGG-IDH1 +0.75% +1.05% +0.27% +2.10% +2.66%
Avg increases +4.12% +1.72% +1.70% +2.68% +0.72%
Table 5: C-Index of Survival Analysis with ABMIL on 9 TCGA datasets. Best performing model for each metric is bolded. 95% CI is included in parentheses.
Task R50 PLIP CONCH UNI mSTAR (Ours)
BRCA (1023) 0.5696 (0.4446,0.6945) 0.5915 (0.4764,0.7065) 0.6940 (0.6052,0.7828) 0.6908 (0.5860,0.7957) 0.7076 (0.6180,0.7972)
CRC (579) 0.5710 (0.4562,0.6857) 0.7076 (0.6186,0.7966) 0.7026 (0.6149,0.7902) 0.6906 (0.6071,0.7742) 0.6895 (0.6059,0.7731)
GBMLGG (830) 0.7810 (0.7425,0.8196) 0.7716 (0.7307,0.8124) 0.7704 (0.7269,0.8139) 0.7905 (0.7471,0.8339) 0.7923 (0.7497,0.8349)
HNSC (441) 0.6087 (0.5297,0.6878) 0.6569 (0.5783,0.7355) 0.6294 (0.5478,0.7109) 0.6516 (0.5718,0.7315) 0.6604 (0.5810,0.7398)
KIRC (498) 0.6258 (0.5392,0.7123) 0.6613 (0.5751,0.7474) 0.6957 (0.6178,0.7736) 0.7155 (0.6496,0.7814) 0.7027 (0.6137,0.7917)
LUAD (455) 0.5995 (0.5071,0.6920) 0.5906 (0.4978,0.6834) 0.6233 (0.5302,0.7163) 0.6312 (0.5316,0.7308) 0.6329 (0.5353,0.7305)
LUSC (452) 0.5305 (0.4395,0.6214) 0.5464 (0.4614,0.6314) 0.6045 (0.5226,0.6865) 0.6273 (0.5502,0.7044) 0.6323 (0.5538,0.7108)
SKCM (415) 0.5811 (0.5074,0.6547) 0.5648 (0.4893,0.6402) 0.6210 (0.5403,0.7017) 0.6254 (0.5478,0.7030) 0.6281 (0.5520,0.7041)
UCEC (495) 0.7450 (0.6630,0.8271) 0.7738 (0.6835,0.8640) 0.8082 (0.7080,0.9085) 0.7845 (0.6833,0.8858) 0.8092 (0.7227,0.8957)
Avg C-Index 0.6236 0.6516 0.6832 0.6897 0.6950
Avg Rank 4.56 3.67 3.00 2.33 1.44
Table 6: C-Index of Survival Analysis with TransMIL on 9 TCGA datasets. Best performing model for each metric is bolded. 95% CI is included in parentheses.
Task R50 PLIP CONCH UNI mSTAR (Ours)
BRCA (1023) 0.6028 (0.4883,0.7174) 0.6193 (0.5436,0.6950) 0.6326 (0.5350,0.7302) 0.6649 (0.5657,0.7641) 0.6937 (0.6098,0.7776)
CRC (579) 0.6222 (0.5148,0.7296) 0.6595 (0.5537,0.7654) 0.6211 (0.5042,0.7380) 0.6277 (0.5074,0.7480) 0.6672 (0.5654,0.7690)
GBMLGG (830) 0.7830 (0.7456,0.8204) 0.7731 (0.7344,0.8117) 0.7778 (0.7384,0.8172) 0.7768 (0.7347,0.8190) 0.7829 (0.7404,0.8254)
HNSC (441) 0.6078 (0.5280,0.6875) 0.6255 (0.5498,0.7011) 0.5783 (0.4995,0.6572) 0.6811 (0.6114,0.7508) 0.6824 (0.6109,0.7539)
KIRC (498) 0.6868 (0.6109,0.7627) 0.6659 (0.5853,0.7465) 0.7080 (0.6327,0.7833) 0.7071 (0.6281,0.7860) 0.7171 (0.6369,0.7972)
LUAD (455) 0.5745 (0.4660,0.6829) 0.5902 (0.4868,0.6936) 0.6121 (0.5100,0.7143) 0.5980 (0.4952,0.7009) 0.6455 (0.5494,0.7417)
LUSC (452) 0.5983 (0.5068,0.6899) 0.5375 (0.4481,0.6270) 0.5651 (0.4732,0.6571) 0.5985 (0.5082,0.6888) 0.6272 (0.5404,0.7140)
SKCM (415) 0.5569 (0.4799,0.6340) 0.5537 (0.4799,0.6276) 0.5941 (0.5218,0.6664) 0.5931 (0.5219,0.6643) 0.6311 (0.5647,0.6976)
UCEC (495) 0.6973 (0.5729,0.8218) 0.7004 (0.6021,0.7988) 0.7651 (0.6766,0.8537) 0.7561 (0.6756,0.8366) 0.7754 (0.6988,0.8520)
Avg C-Index 0.6366 0.6361 0.6505 0.6670 0.6914
Avg Rank 3.89 4.11 3.11 2.78 1.11
Table 7: C-Index of Survival Analysis with the pretrained TransMIL on 9 TCGA datasets. The performance gains of the aggregator (TransMIL) between “before”” and “after” pretraining. ’Positive’(+) values indicate performance increases ($\uparrow$) compared with the ones equipped with TransMIL built from scratch, and vice versa. The best improvement is bolded.
Task R50 PLIP CONCH UNI mSTAR (Ours)
BRCA (1023) -2.06% +1.40% +2.34% +0.08% +0.08%
CRC (579) -9.90% +2.61% +3.17% +3.24% +1.56%
GBMLGG (830) +0.54% +1.93% +0.37% -0.38% -1.11%
HNSC (441) +1.04% +1.69% +5.39% -1.02% +0.21%
KIRC (498) -4.73% +3.57% +2.58% +0.95% +1.69%
LUAD (455) +6.98% +2.04% -0.57% +4.78% +0.15%
LUSC (452) +0.50% +1.51% +2.95% -3.19% -1.41%
SKCM (415) +4.29% +0.55% +2.97% +0.68% +1.53%
UCEC (495) -3.55% +7.79% +2.10% +4.20% +3.87%
Avg increases -0.77% +2.57% +2.37% +1.04% +0.73%
Table 8: C-Index of Multimodal Survival Analysis with MCAT on 9 TCGA datasets. Best performing model for each metric is bolded. 95% CI is included in parentheses.
Task R50 PLIP CONCH UNI mSTAR (Ours)
BRCA 0.7215 (0.6382,0.8048) 0.7304 (0.6428,0.8180) 0.6266 (0.5019,0.7513) 0.6928 (0.6034,0.7822) 0.7489 (0.6703,0.8275)
CRC 0.6996 (0.5994,0.7998) 0.6676 (0.5669,0.7683) 0.6635 (0.5601,0.7669) 0.6481 (0.5474,0.7488) 0.6989 (0.5943,0.8035)
GBMLGG 0.8713 (0.8343,0.9083) 0.8784 (0.8433,0.9135) 0.8886 (0.8553,0.9219) 0.8810 (0.8471,0.9149) 0.8896 (0.8610,0.9182)
HNSC 0.6830 (0.6054,0.7606) 0.6444 (0.5677,0.7211) 0.6646 (0.6008,0.7284) 0.6874 (0.6125,0.7623) 0.6907 (0.6157,0.7657)
KIRC 0.6856 (0.6130,0.7582) 0.6830 (0.6088,0.7572) 0.6996 (0.6335,0.7657) 0.7091 (0.6291,0.7891) 0.7084 (0.6407,0.7761)
LUAD 0.6375 (0.5326,0.7424) 0.6505 (0.5576,0.7434) 0.6628 (0.5681,0.7575) 0.6595 (0.5733,0.7457) 0.6953 (0.6073,0.7833)
LUSC 0.5373 (0.4420,0.6326) 0.5499 (0.4654,0.6344) 0.5526 (0.4660,0.6392) 0.5493 (0.4488,0.6498) 0.5537 (0.4590,0.6484)
SKCM 0.6724 (0.5944,0.7504) 0.6641 (0.5941,0.7341) 0.6522 (0.5781,0.7263) 0.6549 (0.5794,0.7304) 0.6520 (0.5840,0.7200)
UCEC 0.6406 (0.4967,0.7845) 0.6380 (0.5007,0.7753) 0.6754 (0.5532,0.7976) 0.6807 (0.5681,0.7933) 0.6811 (0.5751,0.7871)
Avg C-Index 0.6832 0.6785 0.6762 0.6848 0.7021
Avg Rank 3.44 3.67 3.22 3.00 1.67
Table 9: C-Index of Multimodal Survival Analysis with Porpoise on 9 TCGA datasets. Best performing model for each metric is bolded. 95% CI is included in parentheses.
R50 PLIP CONCH UNI mSTAR (Ours)
BRCA 0.7366 (0.6480,0.8252) 0.7244 (0.6297,0.8191) 0.7259 (0.6369,0.8149) 0.7474 (0.6651,0.8297) 0.7456 (0.6515,0.8397)
CRC 0.6113 (0.4978,0.7248) 0.6639 (0.5571,0.7707) 0.6517 (0.5535,0.7499) 0.6599 (0.5621,0.7577) 0.7020 (0.6085,0.7955)
GBMLGG 0.8805 (0.8472,0.9138) 0.8748 (0.8407,0.9089) 0.8786 (0.8406,0.9166) 0.8948 (0.8627,0.9269) 0.8955 (0.8643,0.9267)
HNSC 0.6982 (0.6271,0.7693) 0.6806 (0.5977,0.7635) 0.6730 (0.5981,0.7479) 0.7111 (0.6415,0.7807) 0.7189 (0.6474,0.7904)
KIRC 0.6681 (0.5934,0.7428) 0.6715 (0.5921,0.7509) 0.6802 (0.5973,0.7631) 0.6767 (0.6046,0.7488) 0.6834 (0.6115,0.7553)
LUAD 0.6656 (0.5792,0.7520) 0.6222 (0.5266,0.7178) 0.6139 (0.5265,0.7013) 0.6295 (0.5454,0.7136) 0.6300 (0.5455,0.7145)
LUSC 0.5364 (0.4490,0.6238) 0.5100 (0.4136,0.6064) 0.4911 (0.3980,0.5842) 0.5794 (0.4859,0.6729) 0.5895 (0.5027,0.6763)
SKCM 0.6383 (0.5666,0.7100) 0.6215 (0.5498,0.6932) 0.6253 (0.5510,0.6996) 0.5995 (0.5233,0.6757) 0.6182 (0.5443,0.6921)
UCEC 0.6154 (0.4809,0.7499) 0.6301 (0.4968,0.7634) 0.6565 (0.5218,0.7912) 0.6565 (0.5385,0.7745) 0.6724 (0.5340,0.8108)
Avg C-Index 0.6723 0.6666 0.6662 0.6839 0.6951
Avg Rank 3.22 3.89 3.72 2.61 1.56
Table 10: C-Index of Multimodal Survival Analysis with MOTCat on 9 TCGA datasets. Best performing model for each metric is bolded. 95% CI is included in parentheses.
R50 PLIP CONCH UNI mSTAR (Ours)
BRCA 0.7030 (0.6075,0.7985) 0.6732 (0.5685,0.7779) 0.6781 (0.5789,0.7773) 0.6970 (0.5814,0.8126) 0.7278 (0.6276,0.8280)
CRC 0.6867 (0.5926,0.7808) 0.6695 (0.5635,0.7755) 0.6643 (0.5553,0.7733) 0.6835 (0.5765,0.7905) 0.7244 (0.6327,0.8161)
GBMLGG 0.8838 (0.8526,0.9150) 0.8736 (0.8375,0.9097) 0.8867 (0.8555,0.9179) 0.8899 (0.8583,0.9215) 0.8889 (0.8583,0.9195)
HNSC 0.6776 (0.6045,0.7507) 0.6731 (0.6023,0.7439) 0.6679 (0.5852,0.7506) 0.6736 (0.5930,0.7542) 0.7256 (0.6546,0.7966)
KIRC 0.7124 (0.6467,0.7781) 0.6890 (0.6108,0.7672) 0.6927 (0.6163,0.7691) 0.7050 (0.6286,0.7814) 0.7085 (0.6403,0.7767)
LUAD 0.6463 (0.5559,0.7367) 0.6472 (0.5549,0.7395) 0.6685 (0.5758,0.7612) 0.6710 (0.5875,0.7545) 0.7004 (0.6138,0.7870)
LUSC 0.5618 (0.4748,0.6488) 0.5564 (0.4621,0.6507) 0.5643 (0.4704,0.6582) 0.5654 (0.4780,0.6528) 0.5855 (0.4957,0.6753)
SKCM 0.6660 (0.5874,0.7446) 0.6607 (0.5868,0.7346) 0.6454 (0.5715,0.7193) 0.6604 (0.5834,0.7374) 0.6715 (0.5972,0.7458)
UCEC 0.6446 (0.5190,0.7702) 0.6899 (0.5601,0.8197) 0.7006 (0.5869,0.8143) 0.7031 (0.5943,0.8119) 0.7353 (0.6453,0.8253)
Avg C-Index 0.6869 0.6814 0.6854 0.6943 0.7187
Avg Rank 3.00 4.33 3.89 2.56 1.22
Table 11: C-Index of Multimodal Survival Analysis with CMTA on 9 TCGA datasets. Best performing model for each metric is bolded. 95% CI is included in parentheses.
R50 PLIP CONCH UNI mSTAR (Ours)
BRCA 0.7118 (0.6224,0.8012) 0.7039 (0.5945,0.8133) 0.7196 (0.6157,0.8235) 0.7339 (0.6351,0.8327) 0.7375 (0.6560,0.8190)
CRC 0.6719 (0.5610,0.7828) 0.6651 (0.5514,0.7788) 0.6580 (0.5469,0.7691) 0.6727 (0.5714,0.7740) 0.7040 (0.6080,0.8000)
GBMLGG 0.8798 (0.8447,0.9149) 0.8820 (0.8479,0.9161) 0.8822 (0.8501,0.9143) 0.9016 (0.8726,0.9306) 0.9051 (0.8798,0.9304)
HNSC 0.6764 (0.6051,0.7477) 0.6782 (0.6059,0.7505) 0.6744 (0.5964,0.7524) 0.6710 (0.5928,0.7492) 0.7092 (0.6430,0.7754)
KIRC 0.6858 (0.6158,0.7558) 0.6752 (0.5984,0.7520) 0.6963 (0.6175,0.7751) 0.7000 (0.6100,0.7900) 0.7065 (0.6279,0.7851)
LUAD 0.6502 (0.5536,0.7468) 0.6301 (0.5386,0.7216) 0.6639 (0.5645,0.7633) 0.6380 (0.5376,0.7384) 0.6574 (0.5684,0.7464)
LUSC 0.5473 (0.4569,0.6377) 0.5234 (0.4303,0.6165) 0.5352 (0.4413,0.6291) 0.5909 (0.4980,0.6838) 0.6320 (0.5497,0.7143)
SKCM 0.6514 (0.5783,0.7245) 0.6546 (0.5786,0.7306) 0.6413 (0.5643,0.7183) 0.6446 (0.5693,0.7199) 0.6488 (0.5731,0.7245)
UCEC 0.6986 (0.5786,0.8186) 0.6676 (0.5504,0.7848) 0.6932 (0.5736,0.8128) 0.7307 (0.6100,0.8514) 0.7555 (0.6669,0.8441)
Avg C-Index 0.6859 0.6756 0.6849 0.6982 0.7173
Avg Rank 3.33 4.00 3.56 2.78 1.33
Table 12: Average Macro-AUC of Few-shot Slide Classification across different shots on 6 datasets. Best performing model for each metric is in bold. Mean±std is presented.
R50 PLIP CONCH UNI mSTAR(Ours)
CAMELYON 0.525±0.019 0.546±0.033 0.504±0.020 0.666±0.087 0.707±0.063
PANDA 0.574±0.020 0.664±0.045 0.648±0.031 0.604±0.022 0.621±0.030
UBC-OCEAN 0.589±0.022 0.723±0.012 0.715±0.027 0.764±0.023 0.774±0.032
BCNB-ER 0.563±0.025 0.582±0.034 0.603±0.068 0.623±0.049 0.624±0.065
BCNB-PR 0.535±0.049 0.599±0.027 0.614±0.017 0.640±0.039 0.660±0.030
BCNB-HER2 0.559±0.054 0.514±0.014 0.524±0.034 0.574±0.039 0.542±0.016
Avg Macro-AUC 0.5575 0.6046 0.6012 0.6457 0.6548
Avg Rank 4.33 3.33 3.50 2.17 1.67
Table 13: Macro-AUC of Few-shot Slide Classification on CAMELYON across different shots. Best performing model for each metric is bolded. Mean±std is presented.
k-shot R50 PLIP CONCH UNI mSTAR(Ours)
k = 1 0.515±0.053 0.511±0.019 0.523±0.037 0.571±0.114 0.659±0.086
k = 2 0.486±0.030 0.521±0.026 0.489±0.045 0.587±0.107 0.600±0.090
k = 4 0.545±0.035 0.517±0.037 0.487±0.067 0.561±0.075 0.660±0.023
k = 8 0.539±0.051 0.537±0.012 0.539±0.061 0.678±0.083 0.742±0.051
k = 16 0.522±0.067 0.571±0.015 0.495±0.053 0.714±0.050 0.737±0.014
k = 32 0.531±0.020 0.551±0.031 0.514±0.094 0.764±0.036 0.788±0.029
k = 64 0.537±0.083 0.612±0.152 0.480±0.136 0.787±0.079 0.766±0.096
Avg Macro-AUC 0.5250 0.5457 0.5039 0.6660 0.7074
Avg Rank 3.93 3.71 4.36 1.86 1.14
Table 14: Macro-AUC of Few-shot Slide Classification on PANDA across different shots. Best performing model for each metric is bolded. Mean±std is presented.
k-shot R50 PLIP CONCH UNI mSTAR(Ours)
k = 1 0.552±0.057 0.564±0.052 0.584±0.064 0.564±0.059 0.570±0.056
k = 2 0.535±0.036 0.619±0.033 0.634±0.037 0.572±0.040 0.574±0.057
k = 4 0.576±0.077 0.677±0.006 0.675±0.028 0.623±0.042 0.632±0.057
k = 8 0.588±0.014 0.678±0.010 0.660±0.031 0.609±0.010 0.643±0.023
k = 16 0.566±0.030 0.678±0.022 0.693±0.018 0.617±0.041 0.629±0.039
k = 32 0.590±0.024 0.700±0.008 0.652±0.026 0.621±0.034 0.624±0.035
k = 64 0.588±0.038 0.696±0.008 0.655±0.026 0.623±0.011 0.650±0.028
k = 128 0.598±0.005 0.701±0.010 0.634±0.022 0.606±0.009 0.646±0.025
Avg Macro-AUC 0.5741 0.6641 0.6484 0.6044 0.6210
Avg Rank 5.00 1.56 1.75 3.94 2.75
Table 15: Macro-AUC of Few-shot Slide Classification on UBC-OCEAN across different shots. Best performing model for each metric is bolded. Mean±std is presented.
k-shot R50 PLIP CONCH UNI mSTAR(Ours)
k = 1 0.585±0.069 0.706±0.036 0.744±0.035 0.726±0.048 0.722±0.049
k = 2 0.625±0.024 0.739±0.033 0.672±0.078 0.786±0.017 0.790±0.019
k = 4 0.584±0.047 0.722±0.061 0.729±0.037 0.763±0.035 0.779±0.034
k = 8 0.563±0.077 0.726±0.071 0.714±0.042 0.779±0.069 0.806±0.089
Avg Macro-AUC 0.5893 0.7232 0.7147 0.7635 0.7742
Avg Rank 5.00 3.50 3.00 2.00 1.50
Table 16: Macro-AUC of Few-shot Slide Classification on BCNB-ER across different shots. Best performing model for each metric is bolded. Mean±std is presented.
k-shot R50 PLIP CONCH UNI mSTAR(Ours)
k = 1 0.514±0.084 0.539±0.104 0.627±0.106 0.595±0.090 0.606±0.112
k = 2 0.598±0.048 0.633±0.061 0.679±0.082 0.614±0.095 0.617±0.143
k = 4 0.571±0.036 0.557±0.079 0.674±0.110 0.549±0.162 0.527±0.169
k = 8 0.559±0.073 0.553±0.123 0.610±0.071 0.670±0.115 0.621±0.116
k = 16 0.572±0.070 0.608±0.048 0.509±0.083 0.614±0.081 0.622±0.085
k = 32 0.561±0.054 0.604±0.069 0.517±0.096 0.697±0.047 0.748±0.051
Avg Macro-AUC 0.5625 0.5823 0.6027 0.6232 0.6235
Avg Rank 4.00 3.33 2.67 2.67 2.33
Table 17: Macro-AUC of Few-shot Slide Classification on BCNB-PR across different shots. Best performing model for each metric is bolded. Mean±std is presented.
k-shot R50 PLIP CONCH UNI mSTAR(Ours)
k = 1 0.551±0.035 0.576±0.069 0.614±0.096 0.640±0.096 0.648±0.137
k = 2 0.507±0.025 0.582±0.037 0.582±0.049 0.595±0.075 0.645±0.071
k = 4 0.532±0.092 0.567±0.093 0.608±0.102 0.606±0.107 0.629±0.160
k = 8 0.602±0.022 0.629±0.071 0.621±0.087 0.656±0.039 0.715±0.028
k = 16 0.571±0.120 0.596±0.058 0.625±0.101 0.716±0.029 0.686±0.057
k = 32 0.447±0.058 0.641±0.010 0.634±0.078 0.645±0.149 0.638±0.073
Avg Macro-AUC 0.5350 0.5985 0.6140 0.6430 0.6602
Avg Rank 5.00 3.42 3.25 1.83 1.50
Table 18: Macro-AUC of Few-shot Slide Classification on BCNB-HER2 across different shots. Best performing model for each metric is bolded. Mean±std is presented.
k-shot R50 PLIP CONCH UNI mSTAR(Ours)
k = 1 0.493±0.050 0.491±0.067 0.502±0.068 0.523±0.080 0.534±0.081
k = 2 0.532±0.027 0.507±0.033 0.548±0.032 0.565±0.079 0.554±0.064
k = 4 0.498±0.051 0.522±0.046 0.562±0.060 0.530±0.037 0.518±0.071
k = 8 0.611±0.033 0.509±0.029 0.555±0.062 0.589±0.053 0.561±0.041
k = 16 0.629±0.037 0.518±0.025 0.506±0.013 0.610±0.080 0.556±0.015
k = 32 0.592±0.056 0.536±0.029 0.468±0.019 0.627±0.040 0.531±0.036
Avg Macro-AUC 0.5592 0.5138 0.5235 0.5740 0.5423
Avg Rank 2.83 4.17 3.50 1.67 2.83
Table 19: Macro-AUC of Zero-shot Slide Classification on 6 datasets. Best performing model for each metric is in bold. 95% CI is included in parentheses.
R50 PLIP CONCH UNI mSTAR(Ours)
CAMELYON 0.541 (0.455,0.627) 0.622 (0.534,0.710) 0.638 (0.550,0.726) 0.685 (0.601,0.769) 0.742 (0.668,0.817)
PANDA 0.480 (0.466,0.494) 0.561 (0.547,0.575) 0.582 (0.566,0.598) 0.450 (0.434,0.466) 0.462 (0.446,0.478)
UBC-OCEAN 0.445 (0.369,0.521) 0.762 (0.691,0.833) 0.370 (0.307,0.433) 0.694 (0.620,0.769) 0.692 (0.621,0.763)
BCNB-ER 0.631 (0.535,0.727) 0.563 (0.473,0.653) 0.500 (0.500,0.500) 0.618 (0.524,0.712) 0.725 (0.641,0.809)
BCNB-PR 0.554 (0.458,0.650) 0.523 (0.435,0.611) 0.500 (0.500,0.500) 0.682 (0.598,0.766) 0.722 (0.644,0.800)
BCNB-HER2 0.549 (0.455,0.643) 0.516 (0.430,0.602) 0.518 (0.428,0.608) 0.465 (0.373,0.557) 0.435 (0.345,0.525)
Avg Macro-AUC 0.5333 0.5912 0.5180 0.5990 0.6297
Avg Rank 3.00 3.00 3.50 3.17 2.33
Table 20: Performance on Pathological Report Generation. Best performing model for each metric is in bold. 6 metrics are reported.
Metric R50 PLIP CONCH UNI mSTAR (Ours)
BLEU-1 0.2510 0.2288 0.2491 0.3590 0.3662
BLEU-2 0.1240 0.1189 0.1527 0.2490 0.2585
BLEU-3 0.0692 0.0727 0.1059 0.1884 0.1981
BLEU-4 0.0410 0.0491 0.0787 0.1490 0.1583
METEOR 0.0898 0.0915 0.1099 0.1570 0.1613
ROUGE-L 0.1825 0.1774 0.2153 0.2991 0.3070
Table 21: Data splits (cases) of 9 cancer datasets on TCGA. CRC=COAD+READ and GBMLGG=GBM+LGG. Apart from these specific cancer types, other TCGA datasets were used for pretraining without data partitioning.
Cancer Type Train Validation Test Total
BRCA 716 102 307 1023
CRC 405 58 116 579
GBMLGG 581 83 166 830
HNSC 308 44 89 441
KIRC 348 50 100 498
LUAD 318 45 92 455
LUSC 316 45 91 452
SKCM 290 41 84 415
UCEC 346 49 100 495
Table 22: Prompts used in GPT-4 for cleaning pathology reports.
Helping me to check the formatting and spelling of the supplied text,
including some incorrect use of punctuation like mis-using of ’X’ and ’x’
and capitalization and deletion of some words of unknown meaning as well.
Table 23: Hyperparameters of ABMIL and TransMIL used for Slide Classification. A single 80GB NVIDIA H800 GPU was used for training.
Hyperparameter Value
Input dim 512/1024
Hidden dim 512
Dropout 0.25
Batch size 1
Epochs 30
Optimizer Adam
Learning rate 2e-4
Scheduler Cosine
Weight decay 1e-5
Table 24: Hyperparameters of ABMIL and TransMIL used for Survival Analysis. A single 80GB NVIDIA H800 GPU was used for training.
Hyperparameter Value
Input dim 512/1024
Hidden dim 512
Dropout 0.25
Batch size 1
Epochs 30
Optimizer Adam
Learning rate 2e-4
Scheduler Cosine
Weight decay 1e-5
Table 25: Architecture Hyperparameters of Porpoise used for Multimodal Survival Analysis. A single 80GB NVIDIA H800 GPU was used for training. $D$ is the number of genes, varying across different datasets.
Hyperparameter WSI RNASeq
Input dim 512/1024 $D$
Hidden dim $512\rightarrow 256$ 256
Dropout 0.1 0.1
Feature dim after fusion 256
Table 26: Architecture Hyperparameters of MCAT and MOTCat used for Multimodal Survival Analysis. A single 80GB NVIDIA H800 GPU was used for training. $D$ is the number of genes, varying across different unique functional categories and different datasets. $256\times 2\rightarrow 256$ refers to 256-dimensional features from two modalities were fused into a 256-dimensional slide-level feature.
Hyperparameter WSI RNASeq
Input dim 512/1024 $6\times D$
Hidden dim 256 $6\times 256$
Dropout 0.25 0.25
Feature dim after fusion $256\times 2\rightarrow 256$
Table 27: Architecture Hyperparameters of CMTA used for Multimodal Survival Analysis. A single 80GB NVIDIA H800 GPU was used for training. $D$ is the number of genes, varying across different datasets. $256\times 2\rightarrow 256$ refers to 256-dimensional features from two modalities were fused into a 256-dimensional slide-level feature.
Hyperparameter WSI RNASeq
Input dim 512/1024 $6\times D$
Hidden dim 256 $6\times(1024\rightarrow 256)$
Dropout 0.25 0.25
Feature dim after fusion $256\times 2\rightarrow 256$
Table 28: Training Hyperparameters used in Multimodal Survival Analysis. A single 80GB NVIDIA H800 GPU was used for training.
Hyperparameter Value
Batch size 1
Epochs 30
Optimizer Adam
Learning rate 2e-4
Scheduler Cosine
Weight decay 1e-5
Table 29: Architecture Hyperparameters of HistGen used for Pathology Report Generation. A single 80GB NVIDIA H800 GPU was used for training. $D$ is the feature dimension of pathological patch, varying across different FMs.
Hyperparameter Value
Input dim $D$
Hidden dim 512
Layers of decoder 3 (8 heads)
Dimension of cross-modal context $512\times 2048$
Optimizer Adam
Learning rate 1e-4
Beam size 3
Table 30: Prompt templates for all tasks that required prompts. In these templates, the placeholder ”CLASSNAME” was replaced with the actual name of the class. The class prompts of CLASSNAME for each task are presented in  31-36.
# Prompt
1 CLASSNAME.
2 a photomicrograph showing CLASSNAME.
3 a photomicrograph of CLASSNAME.
4 an image of CLASSNAME.
5 an image showing CLASSNAME.
6 an example of CLASSNAME.
7 CLASSNAME is shown.
8 this is CLASSNAME.
9 there is CLASSNAME.
10 a histopathological image showing CLASSNAME.
11 a histopathological image of CLASSNAME.
12 a histopathological photograph of CLASSNAME.
13 a histopathological photograph showing CLASSNAME.
14 shows CLASSNAME.
15 presence of CLASSNAME.
16 CLASSNAME is present.
17 an H&E stained image of CLASSNAME.
18 an H&E stained image showing CLASSNAME.
19 an H&E image showing CLASSNAME.
20 an H&E image of CLASSNAME.
21 CLASSNAME, H&E stain.
22 CLASSNAME, H&E.
23 Close-up view of CLASSNAME.
24 Detailed image capturing CLASSNAME.
25 Microscopic analysis reveals CLASSNAME.
26 CLASSNAME under magnification.
27 The structure of CLASSNAME is visible.
28 Diagnosing CLASSNAME.
29 Typical appearance of CLASSNAME.
30 Zoomed image of CLASSNAME.
31 CLASSNAME in detail.
32 CLASSNAME at high magnification.
33 CLASSNAME in a clinical sample.
34 Clinical representation of CLASSNAME.
35 Pathological review of CLASSNAME.
36 Characteristics of CLASSNAME observed.
37 CLASSNAME identified.
38 Diagnosis: CLASSNAME.
39 Specimen showing CLASSNAME.
40 View of CLASSNAME with immunohistochemistry.
41 CLASSNAME with special staining.
42 Observation of CLASSNAME with fluorescent staining.
Table 31: Class prompts for Breast Metastasis Detection.
Task Class CLASSNAME
Breast Metastasis Detection Tumor metastatic cancer
lymph node with metastatic cancer
metastatic breast cancer in lymph node
tumor
Normal non-metastatic
lymph node without metastatic cancer
non-metastatic lymph node
normal
Table 32: Class prompts of PANDA for Prostate ISUP grading.
Task Class CLASSNAME
Prostate ISUP grading 0 benign tissue,
healthy tissue,
non-cancerous tissue,
normal tissue
1 Gleason score 6,
ISUP grade 1,
low-grade cancer,
Gleason pattern 3+3
2 Gleason score 7 (3+4),
ISUP grade 2,
moderately differentiated cancer,
Gleason pattern 3+4
3 Gleason score 7 (4+3),
ISUP grade 3,
moderately differentiated cancer,
Gleason pattern 4+3
4 Gleason score 8,
ISUP grade 4,
high-grade cancer,
Gleason pattern 4+4, 3+5, or 5+3
5 Gleason score 9 or 10,
ISUP grade 5,
high-grade aggressive cancer,
Gleason pattern 4+5, 5+4, or 5+5
Table 33: Class prompts of UBC-OCEAN for Ovarian Cancer Subtyping.
Task Class CLASSNAME
Ovarian Cancer Subtyping HGSC high-grade serous carcinoma,
ovarian high-grade serous carcinoma,
high-grade serous carcinoma of ovary
LGSC low-grade serous carcinoma,
ovarian low-grade serous carcinoma,
low-grade serous carcinoma of ovary
EC endometrioid carcinoma,
ovarian endometrioid carcinoma,
endometrioid carcinoma of ovary
CC clear cell carcinoma,
ovarian clear cell carcinoma,
clear cell carcinoma of ovary
MC mucinous carcinoma,
ovarian mucinous carcinoma,
mucinous carcinoma of ovary
Table 34: Class prompts of BCNB for ER Prediction.
Task Class CLASSNAME
ER Prediction Positive ER positive,
ER+,
ER positive breast cancer,
Estrogen Receptor positive,
ER+ breast cancer,
ER positive carcinoma,
Estrogen Receptor positive breast cancer
Negative ER negative,
ER-,
ER negative breast cancer,
Estrogen Receptor negative,
ER- breast cancer,
ER negative carcinoma,
Estrogen Receptor negative breast cancer
Table 35: Class prompts of BCNB for PR Prediction.
Task Class CLASSNAME
PR Prediction Positive PR positive,
PR+,
PR positive breast cancer,
Progesterone Receptor positive,
PR+ breast cancer,
PR positive carcinoma,
Progesterone Receptor positive breast cancer
Negative PR negative,
PR-,
PR negative breast cancer,
Progesterone Receptor negative,
PR- breast cancer,
PR negative carcinoma,
Progesterone Receptor negative breast cancer
Table 36: Class prompts of BCNB for HER2 Prediction.
Task Class CLASSNAME
HER2 Prediction Positive HER2 positive,
HER2+,
HER2 positive breast cancer,
HER2 overexpression,
HER2+ breast cancer,
HER2 positive carcinoma,
Human Epidermal growth factor Receptor 2 positive,
Human Epidermal growth factor Receptor 2 overexpression
Negative HER2 negative,
HER2-,
HER2 negative breast cancer,
no HER2 overexpression,
HER2- breast cancer,
HER2 negative carcinoma,
Human Epidermal growth factor Receptor 2 negative